### v1

@@ 1 Introduction
Sleep disorders constitute a pervasive global health challenge, severely diminishing cognitive function, physical health, and overall quality of life. Among the most deleterious of these conditions are Insomnia and Obstructive Sleep Apnea (OSA), which increasingly affect the young adult population [1]. Insomnia is characterized by chronic difficulties in sleep initiation or maintenance, leading to daytime fatigue and significant cognitive alterations [2]. The chronic persistence of these disorders exacerbates the risk of severe comorbidities, including cardiovascular disease and neurocognitive degeneration. Consequently, the accurate and timely diagnosis of emerging sleep pathologies is critical. However, traditional diagnostic modalities, particularly the gold standard polysomnography (PSG), are resource intensive, necessitate expert invigilation, and impose significant temporal constraints. These limitations render PSG inaccessible for long duration monitoring and impractical for large scale population screening [3].
@@
To address the scalability limitations of clinical PSG, Machine Learning (ML) has emerged as a promising paradigm for automated disorder screening. Recent advancements in algorithmic efficacy allow for the analysis of complex physiological signals, demographic profiles, and lifestyle factors to predict diagnostic risks with high precision. Ensemble approaches, deep learning neural networks, and hybrid models have demonstrated the capacity to estimate risk accurately [6]. Furthermore, the integration of these models with modern wearable sensors facilitates cost effective, home based monitoring, thereby mitigating the data acquisition bottlenecks inherent in conventional clinical environments [7, 8]. ML based solutions utilizing limited input modalities, such as nocturnal oximetry or photoplethysmography (PPG), have opened new opportunities for risk stratification even when the patient is at home [9]. Previous studies employing Convolutional Neural Networks (CNN) and Recurrent Neural Networks (RNN), such as the DeepSleepNet architecture, have shown success in staging from single channel EEG, while U time approaches have excelled in segmentation tasks [10].
Despite these advancements, existing ML based solutions often struggle with the heterogeneous nature of lifestyle data and severe class imbalance, which can bias predictions toward healthy classes. While recent benchmarks by Rahman et al. [6] and Monowar et al. [7] have achieved high accuracy, there remains a need for a framework that rigorously handles non linear feature interactions and ensures statistical robustness in its predictions.
In this study, we propose a Dual Pipeline Machine Learning Framework designed to maximize diagnostic precision for multi class sleep disorder screening. By integrating rigorous statistical feature engineering with advanced wrapper based selection methods, our approach addresses the limitations of prior works. The primary contributions of this paper are as follows:
@@
    We introduce a novel Dual Pipeline architecture that processes data through two parallel paths: a statistical pipeline focusing on linear separability via Mutual Information and LDA, and a wrapper based pipeline leveraging Boruta and Autoencoders for non linear feature extraction.
@@
    We achieve state of the art classification performance using Extra Trees and K Nearest Neighbors classifiers, outperforming recent 2025 benchmarks on the Sleep Health and Lifestyle dataset.
@@
    We provide a rigorous statistical validation using the Wilcoxon Signed Rank Test to confirm that the performance improvements are statistically significant and not artifacts of random variance.
@@ 2 Literature Review
The early detection of sleep disorders such as Sleep Apnea and Insomnia is critical for mitigating long term cardiovascular and neurocognitive risks. While traditional polysomnography remains the gold standard, its resource intensity has driven a paradigm shift toward Machine Learning (ML) based screening. This review critically analyzes the evolution of these automated approaches, categorizing them into signal based and lifestyle based modalities, to identify the methodological gaps that necessitate the proposed Dual Pipeline Framework.
2.1 Signal Based Detection Methodologies
The majority of existing literature focuses on detecting disorders, particularly Obstructive Sleep Apnea (OSA), using physiological signals like ECG or SpO2. Kristiansen et al. [11] and Bahrami et al. [12] demonstrated that deep learning architectures could achieve accuracies between 85% and 89% using multi signal inputs. Similarly, Serrano Alarcón et al. [13] utilized 1D CNNs on portable monitor signals to detect events with 84.3% accuracy. While effective, these approaches inherently require physical sensors and continuous monitoring equipment, limiting their scalability for mass population screening or non contact scenarios. Furthermore, studies like Khandelwal et al. [14] acknowledge that reducing data dimensionality remains a challenge in these high frequency signal domains. In the context of Insomnia, researchers such as Tripathi et al. [15] and Sharma et al. [16] successfully leveraged Heart Rate Variability (HRV) features, achieving accuracies exceeding 97% with ensemble and KNN classifiers. However, these models rely heavily on the availability of precise physiological recordings, which may not be accessible in resource constrained primary care settings.
2.2 Lifestyle Based Screening and Multi Disorder Classification
To address the scalability limits of sensor based systems, recent research has pivoted toward screening based on lifestyle and demographic data. Hidayat et al. [17] utilized the Sleep Health and Lifestyle dataset to classify multiple disorders, achieving 92% accuracy with Random Forest. However, their study relied on standard feature selection without deeply addressing the non linear interactions between lifestyle factors. Monowar et al. [7] advanced this field by introducing ensemble stacking and SMOTE for class balancing, pushing accuracy to 96.88%. Despite these improvements, standard ensemble methods often fail to distinctively process linear versus non linear feature relationships, potentially capping performance when heterogeneous data types are involved.
@@ 2.3 Research Gaps and Synthesis
Despite the proliferation of ML models for sleep disorder diagnosis, significant methodological gaps remain. First, most studies treat feature engineering as a monolithic process, applying the same transformation pipeline to all features regardless of their statistical distribution or linear separability. Second, while class imbalance is acknowledged, few studies implement hybrid resampling techniques (like SMOTETomek) that simultaneously generate synthetic samples and clean decision boundaries. Finally, many high performing models lack rigorous statistical verification (such as the Wilcoxon Signed Rank Test) to confirm that performance gains are not merely artifacts of random data splitting.
This study addresses these specific gaps by proposing a Dual Pipeline Framework. Unlike prior works that rely on a single processing path, our approach splits features into statistical and wrapper based pipelines to optimally handle both linear and non linear dependencies. By integrating this with hybrid resampling and rigorous statistical validation, we aim to overcome the precision limitations observed in identifying minority disorder classes in previous lifestyle based studies.
@@
Satapathy et al. (2023) [20] RF, KNN, SVM, CNN plus LSTM Sleep EDF Expanded (30 healthy subjects, PSG) CNN+LSTM 87.4%, KNN 83.65%, SVM 76.04 Compared ML and DL classifiers for automated sleep stage classification Limited to healthy subjects, generalizability unverified, small-scale comparisons.
Dritsas et al. (2024) [21] Logistic Regression, SVM (Linear/Polynomial) NHANES dataset SVM Polynomial accuracy 91.44%, F1 91.44 One-vs-All and One-vs-One ML models for sleep disorder prediction None reported.
Kazemi et al. (2025) [22] 1D ViT MTL, 1D CNN, LSTM, ResNet-1D, CNN+LSTM, SVM, RF, kNN UCI Sleep Center PSG (123 subjects; multimodal PPG, RF, RC, RA) Sleep stage accuracy 78%, apnea detection 74%, $\kappa$ = 0.58–0.66 Multitask 1D ViT for simultaneous sleep stage and disorder classification with attention interpretability Small dataset, focused on respiratory disorders, no real-world deployment.
Calderon et al. (2024) [23] Elastic Net, Bayesian Network (CPDAG), SHAP 31,285 U.S. college students (insomnia + seven comorbid disorders) $R^{2}_{\text{train}}$ = 0.44, $R^{2}_{\text{test}}$ = 0.33, RMSE 5.0–5.47; MDD most predictive Combined ML and Bayesian network to explore directional links; depression symptoms presuppose insomnia Cross-sectional, self-report bias, limited generalizability, no external replication.
Jisha et al. (2024) [24] LSTM, IoT UWB Sensor Custom IR UWB dataset (44 subjects; normal to severe apnea) Accuracy 87–92%, Sensitivity 87–90%, Specificity 83–94% Non-contact IoT + LSTM framework for real-time sleep apnea detection Small dataset, high LSTM latency, no external validation, limited hardware calibration.
@@ 3.2 Proposed Framework Workflow
To address the heterogeneous and imbalanced nature of physiological sleep data, we designed a comprehensive Dual-Pipeline Machine Learning Framework. As illustrated in Figure 1, the workflow begins with data acquisition and extensive preprocessing, including categorical encoding and the generation of interaction features to capture non-linear dependencies. A critical component of this architecture is the parallel execution of two distinct feature engineering pipelines: Pipeline 1 (Statistical) applies RobustScaler and Mutual Information, focusing on linear separability via LDA, while Pipeline 2 (Wrapper-based) utilizes MinMaxScaler and Boruta feature selection, leveraging an Autoencoder for non-linear dimensionality reduction. The processed feature sets converge at the classification stage, where nine supervised algorithms, including advanced ensemble methods, are trained and optimized via Stratified 8-Fold Cross-Validation. This structural design enables a rigorous comparative analysis to identify the optimal preprocessing strategy for minimizing false negatives in sleep disorder screening.
@@ 3.3 Data Preprocessing and Feature Engineering
  • Feature Engineering: To capture non-linear physiological dependencies that single variables may fail to represent, we engineered a set of interaction features. These derived metrics, detailed in Table 5, quantify relationships such as the ratio of sleep duration to physiological stress markers.
@@
Following feature engineering, we conducted a feature importance analysis using tree-based estimators to rank predictors by their contribution to information gain [30, 31]. This analysis, visualized in Figure 2, guided the subsequent selection processes in our dual-pipeline framework.
@@ 3.5 Dual-Pipeline Feature Engineering Architecture
We implemented a dual-pipeline framework to systematically compare traditional statistical transformations against advanced non-linear engineering techniques [29]. This architecture enables the extraction of complementary physiological information through distinct normalization, feature selection, and dimensionality reduction strategies.
@@ 3.5.2 Pipeline 2: Wrapper-based Engineering (Non-Linear Focus)
  • Dimensionality Reduction: Dimensionality reduction is achieved via an Autoencoder. This unsupervised neural network learns a compressed latent representation of the input data. By minimizing reconstruction error, the Autoencoder captures non-linear latent structures that linear methods may miss [39, 40].
@@ 3.6 Machine Learning Classifiers
To ensure a rigorous evaluation of the proposed dual-pipeline framework, we employed a diverse suite of nine supervised learning algorithms, categorized into linear, non-linear, and ensemble paradigms. This selection allows for a comprehensive analysis of how different inductive biases interact with the engineered feature sets.
Linear and Non-Linear Estimators: We utilized Logistic Regression as a baseline linear classifier [41] due to its interpretability and efficacy in high-dimensional spaces where linear separability is assumed. To capture local manifold structures within the physiological data, we employed K-Nearest Neighbors (KNN) [42], a non-parametric method that classifies instances based on feature similarity. Furthermore, to address complex non-linear dependencies without manual feature interaction, we implemented a Multi-Layer Perceptron (MLP) [40, 43]. This neural network architecture leverages hidden layers and non-linear activation functions to approximate complex decision boundaries that linear models may fail to resolve.
@@
Hyperparameter Optimization Strategy: To optimize model generalization, we utilized RandomizedSearchCV rather than an exhaustive grid search [48]. This approach samples a fixed number of parameter settings from specified distributions, offering a superior trade-off between computational efficiency and optimization performance in high-dimensional spaces. The validation process was conducted using Stratified 8-Fold Cross-Validation [49]. While 5-fold or 10-fold schemes are standard, we selected $k=8$ based on the dataset size ($N=374$). This configuration assigns approximately 47 samples to each validation fold. This split provides a statistically significant representation of minority classes (Sleep Apnea and Insomnia) in the validation set, which might be diluted in a 10-fold split, while retaining a larger portion of data for training compared to a 5-fold split, thus ensuring the model learns robust patterns from the limited data available.
@@ 4 Results and Discussion
The empirical evaluation of the proposed framework is divided into two primary phases. The first phase establishes a baseline by assessing the performance of machine learning models on the raw dataset. The second phase analyzes the efficacy of the Dual-Pipeline Framework, highlighting the performance gains achieved through specific combinations of feature engineering and hybrid resampling. A detailed decomposition of these components, including the impact of SMOTETomek and feature selection dynamics, is reserved for the subsequent ablation study.
@@ 4.1 Baseline Model Performance
The results indicate that ensemble and linear models possess inherent predictive capabilities even in the absence of optimization. Specifically, Logistic Regression, Gradient Boosting, and LightGBM demonstrated superior baseline performance, each achieving an accuracy of 97.333%. This suggests that a significant portion of the variance in the dataset is linearly separable. Conversely, the Multi-Layer Perceptron (MLP) Classifier exhibited poor performance with an accuracy of 61.333% and an F1 Score of 32.479%. This suboptimal result for the MLP is attributed to the lack of feature normalization in the raw data, which disproportionately affects neural network convergence. These baseline metrics serve as a critical benchmark for evaluating the relative improvements offered by the proposed dual-pipeline architecture.

### v2

@@ 1 Introduction
Sleep disorders constitute a pervasive global health challenge, severely diminishing cognitive function, physical health, and overall quality of life. Among the most deleterious of these conditions are Insomnia and Obstructive Sleep Apnea (OSA), which increasingly affect the young adult population [1]. Insomnia is characterized by chronic difficulties in sleep initiation or maintenance, leading to daytime fatigue and significant cognitive alterations [2]. The chronic persistence of these disorders exacerbates the risk of severe comorbidities, including cardiovascular disease and neurocognitive degeneration. Consequently, the accurate and timely diagnosis of emerging sleep pathologies is critical. However, traditional diagnostic modalities, particularly the gold standard polysomnography (PSG), are resource-intensive, necessitate expert invigilation, and impose significant temporal constraints. These limitations render PSG inaccessible for long-duration monitoring and impractical for large-scale population screening [3].
@@
To address the scalability limitations of clinical PSG, Machine Learning (ML) has emerged as a promising paradigm for automated disorder screening. Recent advancements in algorithmic efficacy allow for the analysis of complex physiological signals, demographic profiles, and lifestyle factors to predict diagnostic risks with high precision. Ensemble approaches, deep learning neural networks, and hybrid models have demonstrated the capacity to estimate risk accurately [6]. Furthermore, the integration of these models with modern wearable sensors facilitates cost-effective, home-based monitoring, thereby mitigating the data acquisition bottlenecks inherent in conventional clinical environments [7, 8]. ML-based solutions utilizing limited input modalities, such as nocturnal oximetry or photoplethysmography (PPG), have opened new opportunities for risk stratification even when the patient is at home [9]. Previous studies employing Convolutional Neural Networks (CNN) and Recurrent Neural Networks (RNN), such as the DeepSleepNet architecture, have shown success in staging from single-channel Electroencephalogram (EEG)[10].
Despite these advancements, existing ML-based solutions often struggle with the heterogeneous nature of lifestyle data and severe class imbalance, which can bias predictions toward healthy classes. While recent benchmarks by Rahman et al. [6] and Monowar et al. [7] have achieved high accuracy, there remains a need for a framework that rigorously handles nonlinear feature interactions and ensures statistical robustness in its predictions.
In this study, we propose a Dual Pipeline Machine Learning Framework (DPMLF) designed to maximize diagnostic precision for multiclass sleep disorder screening. By integrating rigorous statistical feature engineering with advanced wrapper-based selection methods, our approach addresses the limitations of prior works. The primary contributions of this paper are as follows:
@@
    We introduce a novel Dual Pipeline architecture that processes data through two parallel paths: a statistical pipeline focusing on linear separability via Mutual Information (MI) and Linear Discriminant Analysis (LDA), and a wrapper-based pipeline leveraging Boruta and Autoencoders for nonlinear feature extraction.
@@
    We achieve state-of-the-art classification performance using Extra Trees (Extremely Randomized Trees) and K-Nearest Neighbors (KNN) classifiers, outperforming recent 2025 benchmarks on the Sleep Health and Lifestyle dataset.
@@
    We provide a rigorous statistical validation using the Wilcoxon Signed-Rank Test to confirm that the performance improvements are statistically significant and not artifacts of random variance.
@@ 2 Literature Review
The early detection of sleep disorders such as Sleep Apnea and Insomnia is critical for mitigating long-term cardiovascular and neurocognitive risks. While traditional polysomnography remains the gold standard, its resource intensity has driven a paradigm shift toward ML-based screening. This review critically analyzes the evolution of these automated approaches, categorizing them into signal-based and lifestyle-based modalities, to identify the methodological gaps that necessitate the proposed Dual Pipeline Framework.
2.1 Signal-Based Detection Methodologies
The majority of existing literature focuses on detecting disorders, particularly Obstructive Sleep Apnea (OSA), using physiological signals like ECG or SpO2. Kristiansen et al. [11], and Bahrami et al. [12] demonstrated that deep learning architectures could achieve accuracies between 85% and 89% using multi-signal inputs. Similarly, Serrano Alarcón et al. [13] utilized 1D CNNs on portable monitor signals to detect events with 84.3% accuracy. While effective, these approaches inherently require physical sensors and continuous monitoring equipment, limiting their scalability for mass population screening or non-contact scenarios. Furthermore, studies like Khandelwal et al. [14] acknowledge that reducing data dimensionality remains a challenge in these high-frequency signal domains. In the context of Insomnia, researchers such as Tripathi et al. [15] and Sharma et al. [16] successfully leveraged Heart Rate Variability (HRV) features, achieving accuracies exceeding 97% with ensemble and KNN classifiers. However, these models rely heavily on the availability of precise physiological recordings, which may not be accessible in resource-constrained primary care settings.
2.2 Lifestyle-Based Screening and Multi-Disorder Classification
To address the scalability limits of sensor-based systems, recent research has pivoted toward screening based on lifestyle and demographic data. Hidayat et al. [17] utilized the Sleep Health and Lifestyle dataset to classify multiple disorders, achieving 92% accuracy with Random Forest. However, their study relied on standard feature selection without deeply addressing the nonlinear interactions between lifestyle factors. Monowar et al. [7] advanced this field by introducing ensemble stacking and SMOTE for class balancing, pushing accuracy to 96.88%. Despite these improvements, standard ensemble methods often fail to distinctively process linear versus nonlinear feature relationships, potentially capping performance when heterogeneous data types are involved.
@@ 2.3 Research Gaps and Synthesis
Despite the proliferation of ML models for sleep disorder diagnosis, significant methodological gaps remain. First, most studies treat feature engineering as a monolithic process, applying the same transformation pipeline to all features regardless of their statistical distribution or linear separability. Second, while class imbalance is acknowledged, few studies implement hybrid resampling techniques (like SMOTETomek) that simultaneously generate synthetic samples and clean decision boundaries. Finally, many high-performing models lack rigorous statistical verification (such as the Wilcoxon Signed-Rank Test) to confirm that performance gains are not merely artifacts of random data splitting.
This study addresses these specific gaps by proposing a Dual Pipeline Framework. Unlike prior works that rely on a single processing path, our approach splits features into statistical and wrapper-based pipelines to optimally handle both linear and nonlinear dependencies. By integrating this with hybrid resampling and rigorous statistical validation, we aim to overcome the precision limitations observed in identifying minority disorder classes in previous lifestyle-based studies.
@@
Satapathy et al. (2023) [20] RF, KNN, SVM, CNN plus LSTM Sleep-EDF Dataset CNN+LSTM 87.4%, KNN 83.65%, SVM 76.04%, RF 74.07% Compared ML and DL classifiers for automated sleep stage classification Limited to healthy subjects, generalizability unverified, small-scale comparisons.
Dritsas et al. (2024) [21] Logistic Regression, SVM (Linear/Polynomial) NHANES dataset SVM Polynomial accuracy 91.44%, f-measure 0.914 One-vs-All and One-vs-One ML models for sleep disorder prediction The dataset is small and lacks detailed clinical information on subjects’ profiles
Kazemi et al. (2025) [22] 1D-Vision Transformer(1D-ViT, multitask explainable architecture) Polysomno- graphy(PSG) recordings from UCI Sleep Center (123 subjects) Sleep stage classification accuracy 78%(Cohen’s Kappa  0.66), Sleep apnea classification 74%(Cohen’s Kappa  0.58) Multimodal multitask 1D ViT for simultaneous sleep stage and disorder classification with attention interpretability Small dataset, focused on respiratory disorders, no real-world deployment.
Calderon et al. (2024) [23] Elastic Net, Bayesian Network (CPDAG), SHAP 31,285 U.S. college students (insomnia + seven comorbid disorders) $R^{2}_{\text{train}}$ = 0.44 with RMSE = 5.0, $R^{2}_{\text{test}}$ = 0.33 with RMSE = 5.47; MDD most predictive Combined ML and Bayesian network to explore connections of key disorders with insomnia Cross-sectional, self-report bias, limited generalizability, no external replication.
Jisha et al. (2024) [24] LSTM, IoT UWB Sensor Custom IR UWB dataset Accuracy 87–92%, Sensitivity 87–90%, Specificity 83–94% Non-contact IoT + LSTM framework for real-time sleep apnea detection Small dataset, high LSTM latency, no external validation, limited hardware calibration.
@@ 3.2 Proposed Framework Workflow
To address the heterogeneous and imbalanced nature of physiological sleep data, we designed a comprehensive Dual Pipeline Machine Learning Framework. As illustrated in Figure 1, the workflow begins with data acquisition and extensive preprocessing, including categorical encoding and the generation of interaction features to capture non-linear dependencies. A critical component of this architecture is the parallel execution of two distinct feature engineering pipelines: Pipeline 1 (Statistical) applies RobustScaler and Mutual Information, focusing on linear separability via LDA, while Pipeline 2 (Wrapper-based) utilizes MinMaxScaler and Boruta feature selection, leveraging an Autoencoder for nonlinear dimensionality reduction. The processed feature sets converge at the classification stage, where nine supervised algorithms, including advanced ensemble methods, are trained and optimized via Stratified 8-Fold Cross-Validation. This structural design enables a rigorous comparative analysis to identify the optimal preprocessing strategy for minimizing false negatives in sleep disorder screening.
@@ 3.3 Data Preprocessing and Feature Engineering
  • Feature Engineering: To capture nonlinear physiological dependencies that single variables may fail to represent, we engineered a set of interaction features. These derived metrics, detailed in Table 5, quantify relationships such as the ratio of sleep duration to physiological stress markers.
@@
Following feature engineering, we conducted a feature importance analysis using tree-based estimators to rank predictors by their contribution to information gain [30, 31]. This analysis, visualized in Figure 2, guided the subsequent selection processes in our dual pipeline framework.
@@ 3.5 Dual-Pipeline Feature Engineering Architecture
We implemented a dual-pipeline framework to systematically compare traditional statistical transformations against advanced nonlinear engineering techniques [29]. This architecture enables the extraction of complementary physiological information through distinct normalization, feature selection, and dimensionality reduction strategies.
@@ 3.5.2 Pipeline 2: Wrapper-based Engineering (Non-Linear Focus)
  • Dimensionality Reduction: Dimensionality reduction is achieved via an Autoencoder. This unsupervised neural network learns a compressed latent representation of the input data. By minimizing reconstruction error, the Autoencoder captures nonlinear latent structures that linear methods may miss [39, 40].
@@ 3.6 Machine Learning Classifiers
To ensure a rigorous evaluation of the proposed dual pipeline framework, we employed a diverse suite of nine supervised learning algorithms, categorized into linear, non-linear, and ensemble paradigms. This selection allows for a comprehensive analysis of how different inductive biases interact with the engineered feature sets.
Linear and Non-Linear Estimators: We utilized Logistic Regression as a baseline linear classifier [41] due to its interpretability and efficacy in high-dimensional spaces where linear separability is assumed. To capture local manifold structures within the physiological data, we employed K-Nearest Neighbors (KNN) [42], a non-parametric method that classifies instances based on feature similarity. Furthermore, to address complex non-linear dependencies without manual feature interaction, we implemented a Multi-Layer Perceptron (MLP) [40, 43]. This neural network architecture leverages hidden layers and nonlinear activation functions to approximate complex decision boundaries that linear models may fail to resolve.
@@
Hyperparameter Optimization Strategy: Both GridSearchCV and RandomizedSearchCV were explored during hyperparameter tuning. RandomizedSearchCV allows efficient sampling of large hyperparameter spaces with lower computational cost, while GridSearchCV evaluates all possible parameter combinations, which can be computationally expensive and time-consuming. Therefore, to optimize model generalization, we utilized RandomizedSearchCV rather than an exhaustive grid search [48]. This approach samples a fixed number of parameter settings from specified distributions, offering a superior trade-off between computational efficiency and optimization performance in high-dimensional spaces. The validation process was conducted using Stratified 8-Fold Cross-Validation [49]. While 5-fold or 10-fold schemes are standard, we selected $k=8$ based on the dataset size ($N=374$). This configuration assigns approximately 47 samples to each validation fold. This split provides a statistically significant representation of minority classes (Sleep Apnea and Insomnia) in the validation set, which might be diluted in a 10-fold split, while retaining a larger portion of data for training compared to a 5-fold split, thus ensuring the model learns robust patterns from the limited data available.
@@ 4 Results and Discussion
The empirical evaluation of the proposed framework is divided into two primary phases. The first phase establishes a baseline by assessing the performance of machine learning models on the raw dataset. The second phase analyzes the efficacy of the Dual Pipeline Framework, highlighting the performance gains achieved through specific combinations of feature engineering and hybrid resampling. A detailed decomposition of these components, including the impact of SMOTETomek and feature selection dynamics, is reserved for the subsequent ablation study.
@@ 4.1 Baseline Model Performance
The results indicate that ensemble and linear models possess inherent predictive capabilities even in the absence of optimization. Specifically, Logistic Regression, Gradient Boosting, and LightGBM demonstrated superior baseline performance, each achieving an accuracy of 97.333%. This suggests that a significant portion of the variance in the dataset is linearly separable. Conversely, the Multi-Layer Perceptron (MLP) Classifier exhibited poor performance with an Accuracy of 61.333% and an F1 Score of 32.479%. This suboptimal result for the MLP is attributed to the lack of feature normalization in the raw data, which disproportionately affects neural network convergence. These baseline metrics serve as a critical benchmark for evaluating the relative improvements offered by the proposed dual pipeline architecture.